OpenStudy (anonymous):

Find the coefficient of a^5b^7 in the expansion of (a + b)^12.

OpenStudy (anonymous):

so general formula for comuting (x+y)^n is : \[(x+y)^{n}=\sum_{j=0}^{n}\left(\begin{matrix}n \\ j\end{matrix}\right)x^{n-j}y^{j}\] so you need the one in front of a^5b^7 so that would be substituting for n=12,j=7,x=a , y=b

OpenStudy (anonymous):

We werent really taught that formula so =/ ...is there another way of going about it?

OpenStudy (anonymous):

hm how do you solve it then ?

OpenStudy (anonymous):

what formulas u have

OpenStudy (anonymous):

I've never come across such a question hence putting it on here...but I usually use the simple finding of the factorials then later on multiplying whilst finding the value of a certain coefficient

OpenStudy (anonymous):

ok wait when u look at that formula seems complicated thats just general formula for expansion lets find the coeficient

OpenStudy (anonymous):

Okay, but like since there are no values for a and b I can't do it my way. That formula helps you find it ?

OpenStudy (anonymous):

yeah we should be able to find it so that formula gives us every term in the expansion we want just one so we take just whats inside the sumation so we would have : \[\left(\begin{matrix}n \\ j\end{matrix}\right)x^{n-j}y^{j}\]

OpenStudy (anonymous):

so pluging in our values we would have \[\left(\begin{matrix}12 \\ 7\end{matrix}\right)a^{5}b^{7}\]

OpenStudy (anonymous):

so what we left with is only the combinations \[\left(\begin{matrix}12 \\ 7\end{matrix}\right)\] calculating that we will have the coeficient in front of that term

OpenStudy (anonymous):

do you know to calculate that ?

OpenStudy (anonymous):

Ohhh yeah I can do that. Thank youu

OpenStudy (anonymous):

you are welcomed :)

OpenStudy (anonymous):

But wait

OpenStudy (anonymous):

Ohh nvm 792 got it

OpenStudy (anonymous):

yeap thats correct :)